# Performance Trade-off and Joint Waveform Design for MIMO-OFDM DFRC Systems

### v1

@@ Abstract
Dual-functional radar-communication (DFRC) has attracted considerable attention. This paper considers the frequency-selective multipath fading environment and proposes DFRC waveform design strategies based on multiple-input and multiple-output (MIMO) and orthogonal frequency division multiplexing (OFDM) techniques. In the proposed waveform design strategies, the Cramer-Rao bound (CRB) of the radar system, the inter-stream interference (ISI) and the achievable rate of the communication system, are respectively considered as the performance metrics. In this paper, we focus on the performance trade-off between the radar system and the communication system, and the optimization problems are formulated. In the ISI minimization based waveform design strategy, the optimization problem is convex and can be easily solved. In the achievable rate maximization based waveform design strategy, we propose a water-filling (WF) and sequential quadratic programming (SQP) based algorithm to derive the covariance matrix and the precoding matrix. Simulation results validate the proposed DFRC waveform designs and show that the achievable rate maximization based strategy has a better performance than the ISI minimization based strategy.
@@ I Introduction
The waveform design, which is one of the most challenging researches in the JRC system, is mainly classified into two categories. The first category is that the radar system and the communication system co-exist [10][11], where one system treats the other system as interference [12]. Therefore, multiplexing techniques including space division multiplexing (SDM) [13, 14, 15], time division multiplexing (TDM) [16, 17, 18], frequency division multiplexing (FDM) [19][20], and code division multiplexing (CDM) [21][22] have been applied to mitigate the interference in the co-existing system. However, in these multiplexing schemes, the co-existing waveform has a low utilization because the radar system and the communication system cannot employ the same resources, and two different kinds of waveforms are usually used respectively for radar and communication.
The second category is that the radar system and the communication system are co-designed, which aims to develop a single waveform that can perform radar detection and communication functions simultaneously [23], and this waveform is called dual-functional radar-communication (DFRC) waveform. Compared with the co-existing waveform, the DFRC waveform has a higher degree of integration, thus it is more attractive to both academia and industry [24].
@@
All the above works adopted the single-carrier modulation, and the flat fading channel model was used. However, the frequency-selective multipath fading channel is more reasonable in the high-rate transmission. Orthogonal frequency division multiplexing (OFDM) [36][37] is one of the key technologies in the fifth generation (5G) mobile communication for its robustness to the multipath fading. Besides, it has been proven to be valid for the radar detection in [28]. Xu et al. in [38] presented a complete processing strategy for the measurement of range, velocity, and angle based on OFDM signals. Meanwhile, a trade-off scheme that employed dedicated subcarriers to balance the radar performance and the communication performance was also proposed. However, different subcarriers in [38] used the same precoding matrix, which led to a performance degradation in the DFRC system. The beam strength radiated towards the detection target is usually greater than that towards the communication user. Therefore, a subcarrier power allocation scheme was proposed in [39] to satisfy the different power requirements of the radar system and the communication system. Hu et al. in [40] extended the work of [23] and the PAPR of the designed OFDM waveform was optimized to achieve a better communication performance.
In this paper, we consider a single communication user MIMO system in the frequency-selective multipath fading environment. The DFRC base station (BS) utilizes a single waveform to realize the downlink communication and the target detection simultaneously. Because the MUI minimization strategy used in [23] cannot be directly used in the frequency-selective multipath fading environment, we propose an OFDM based waveform design strategy, which can effectively combat the multipath fading, to minimize the inter-stream interference (ISI) of the communication system and the CRB of the radar system. Furthermore, the achievable rate maximization based waveform design strategy and the joint water-filling and sequential quadratic programming (WF-SQP) algorithm are proposed to further improve the communication performance through the cooperation of different antennas.
@@
1. We focus on the performance trade-off between the communication system and the radar system, and propose two OFDM based DFRC waveform design strategies to combat the frequency-selective multipath fading. In the first proposed DFRC waveform design strategy, we minimize the ISI of the communication system and the CRB of the radar system. The algorithms for designing the transmit covariance matrix and the precoding matrix design algorithms are proposed.
2. Different from the first DFRC waveform design strategy, the achievable rate of the communication system is employed in the second DFRC waveform design strategy. The performance trade-off is also considered. To derive the optimal transmit covariance matrix, a joint water-filling and SQP algorithm is proposed. Furthermore, the precoding matrix is derived analytically.
3. Simulation results are provided and analyzed in detail. Simulation results show that the proposed DFRC waveform design strategies can achieve satisfied communication and radar performances. The second proposed DFRC waveform design strategy achieves better radar and communication performances than the first one.
The rest of this paper is organized as follows. Section II presents the MIMO-OFDM based system model. In Section III, we propose the first OFDM based DFRC waveform design strategy with the aim of minimizing the ISI of the communication and the CRB of the radar. In Section IV, we maximize the achievable rate of the communication system, and propose the second waveform design strategy. Simulation results are provided in Section V. Finally, the conclusion is drawn in Section VI.
@@ II System model
The transmit signal vector $\boldsymbol{g}(n)$ is expressed as
$\boldsymbol{g}(n)={[{g_{1}}(n),{g_{2}}(n),\cdots,{g_{{N_{T}}}}(n)]^{T}},n=1,2,\ldots,{N+N_{CP}-1},$ (1)
where $N$ is the size of inverse discrete Fourier transform (IDFT), $N_{CP}$ is the length of the OFDM cyclic prefix, and $g_{i}(n)$ is the time-domain OFDM signal (including cyclic prefix) transmitted through the $i$th antenna. The transmit signal, $g_{i}(n)$ is given by
$[{g_{i}}(0),{g_{i}}(1),\cdots,{g_{i}}({N_{CP}}-1),{g_{i}}({N_{CP}}),\cdots,{g_{i}}(N+{N_{CP}}-1)]^{T}$ (2)
$=[{{x_{i}}(N-{N_{CP}}),\cdots,{x_{i}}(N-1),{x_{i}}(0),\cdots,{x_{i}}(N-1)}],$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad n=1,2,\ldots,{N+N_{CP}-1},$
where
@@
The frequency-domain signal vector $\boldsymbol{X}(k)$ and the time-domain signal vector $\boldsymbol{x}(n)$ are expressed as
$\boldsymbol{X}(k)={\left[{{X_{1}}(k),{X_{2}}(k),\cdots,{X_{{N_{T}}}}(k)}\right]^{T}},$ $k=0,1,\ldots,N-1,$ (4a)
$\boldsymbol{x}(n)={[{x_{1}}(n),{x_{2}}(n),\cdots,{x_{{N_{T}}}}(n)]^{T}},$ $n=0,1,\ldots,N-1.$ (4b)
$\boldsymbol{X}(k)=\boldsymbol{W}(k)$ $\boldsymbol{S}(k),$ (4c)
@@
$\boldsymbol{R}(i,s)$ $={1\over{{N^{2}}}}E\left\{{\sum\limits_{{k_{1}}}^{N-1}{\sum\limits_{{k_{2}}}^{N-1}{{X_{i}}({k_{1}})}}X_{s}^{*}({k_{2}}){e^{j{{2\pi}\over N}n({k_{1}}-{k_{2}})}}}\right\},$ (5b)
where $E\{\cdot\}$ denotes the expectation of a random variable, $(\cdot)^{H}$ is the conjugate transpose operator, $(\cdot)^{*}$ is the conjugate operator and $\boldsymbol{R}(i,s)$ is the $(i,s)$th entry of $\boldsymbol{R}$. It is supposed that each antenna employs all the subcarriers, and $E\{{\boldsymbol{X}_{i}}({k_{1}})\boldsymbol{X}_{s}^{*}({k_{2}})\}=0$ holds when $k_{1}\neq k_{2}$. Therefore, Eq. (5a) can be rewritten as
@@
$\boldsymbol{R}_{f}(k)=E\{\boldsymbol{X}(k){\boldsymbol{X}^{*}}(k)\}=\boldsymbol{W}(k){\boldsymbol{P}_{S}}(k){\boldsymbol{W}^{H}}(k).$ (7)
@@ II-A MIMO-OFDM Communication Model
${y_{s}}(n)=\sum\limits_{i=1}^{{N_{T}}}{{h_{s,i}}(n)\otimes{x_{i}}(n)}+{z_{s}}(n),s=1,2,\ldots,{N_{R}},$ (9)
where $\otimes$ denotes the convolution operation, and ${z_{s}}(n)$ is the noise component. In the frequency-domain, the received signal vector at the $k$th subcarrier can be expressed as
$\boldsymbol{Y}$ $(k)=\boldsymbol{H}(k)\boldsymbol{X}(k)+\boldsymbol{Z}(k)$ (10)
@@
where ${\boldsymbol{H}}(k)\in{\mathbb{C}^{{N_{R}}\times{N_{T}}}}$ is the frequency-domain channel matrix corresponding to the $k$th subcarrier, ${\boldsymbol{Z}}(k)\in{\mathbb{C}^{N_{R}}}$ is the frequency-domain Guassian noise vector with covariance matrix $\sigma_{\boldsymbol{Z}}^{2}{{\boldsymbol{I}}_{{N_{R}}}}$ , and ${\boldsymbol{I}}_{{N_{R}}}$ is an identical matrix with rank $N_{R}$.
@@
${C_{s}}=\sum\limits_{k=0}^{N-1}\sum\limits_{i=1}^{{N_{R}}}{{{\Delta f}\over B}}{{\log}_{2}}\left(\!{1\!+\!{{E\{{{{\left\|{{\boldsymbol{S}_{i}}(k)}\right\|}^{2}}}\}}\over{E\{{{{\left\|{{\boldsymbol{H}_{i}}(k)\boldsymbol{W}(k)\boldsymbol{S}(k)-{\boldsymbol{S}_{i}}(k)}\right\|}^{2}}}\}+\sigma_{\boldsymbol{Z}}^{2}}}}\!\right),$ (13)
where ${\Delta f}$ is the subcarrier spacing, $B$ is the bandwidth of the OFDM system, ${\boldsymbol{S}_{i}}(k)$ is the $i$th entry of ${\boldsymbol{S}}(k)$, and ${\boldsymbol{H}_{i}}(k)$ is the $i$th row of ${\boldsymbol{H}}(k)$. Note that the effect of the cyclic prefix to the data rate is ignored.
@@
${C_{t}}={\rm{max}}\bigg{\{}\sum\limits_{k=0}^{N-1}{{{\Delta f}\over B}}{{\log}_{2}}\left({\det\left({{\boldsymbol{I}_{{N_{R}}}}+{{\boldsymbol{H}(k){\boldsymbol{R}_{f}}(k){\boldsymbol{H}^{H}}(k)}\over{\sigma_{\boldsymbol{Z}}^{2}}}}\right)}\right)\bigg{\}},$ (14)
@@
Although we focus on the single user communication scenario in this paper, the system model and the following proposed scheme can still be applied in the multiuser scenario, and time division multiple access (TDMA) technique can be used.
@@ II-B MIMO Radar Model
where $\zeta$ is the reflection coefficient and contains the information of the radar cross section (RCS) of the target, ${\boldsymbol{z}_{r}}(n)$ is the Guassian noise vector, $\boldsymbol{A}(\theta)=\boldsymbol{a}(\theta)\boldsymbol{a}(\theta)^{H}$, $\theta$ is the angle of arrival of the target, and the steering vector $\boldsymbol{a}(\theta)$ is given by [33]
@@
where the center of the ULA antennas is chosen as the reference point, and $\Delta$ is the normalized antenna spacing.
The CRB of $\theta$ can be derived as [42]
@@
where $\boldsymbol{\dot{A}}(\theta)={{\partial\boldsymbol{A}(\theta)}\over{\partial\theta}}$, and $\rm{tr}(\cdot)$ is the operator of matrix trace. Substituding $\boldsymbol{A}(\theta)=\boldsymbol{a}(\theta)\boldsymbol{a}(\theta)^{H}$ into Eq. (17), the CRB of $\theta$ can be rewritten as
@@ III Joint waveform design based on CRB and ISI minimization
In this section, we first design the time-domain covariance matrix to minimize the CRB of $\theta$. After that, an optimization problem of minimizing the interference among different information streams for a given time-domain covariance matrix is formulated, and the frequency-domain covariance matrix and the corresponding precoding matrix are designed. Besides, we propose the joint waveform design strategy based on the CRB and ISI minimization to achieve the performance trade-off.
@@ III-A Radar-only problem formulation
[…]
@@
where ${P_{T}}$ is the total transmit power. The optimization problem OP1 is convex [43] and can be solved by CVX tools. The optimal time-domain covariance matrix is denoted by $\boldsymbol{R}_{d}$.

### v2

# Joint Waveform Design for MIMO-OFDM DFRC Systems

@@ Abstract
Dual-functional radar-communication (DFRC) has attracted considerable attention. This paper considers the frequency-selective multipath fading environment and proposes DFRC waveform design schemes based on multiple-input and multiple-output (MIMO) and orthogonal frequency division multiplexing (OFDM) techniques. In the proposed waveform design schemes, the Cramer-Rao bound (CRB) of the radar system, the inter-stream interference (ISI) and the achievable rate of the communication system, are respectively considered as the performance metrics. In this paper, we focus on the performance trade-off between the radar system and the communication system, and the optimization problems are formulated. In the ISI minimization based waveform design scheme, the optimization problem is convex and can be easily solved. In the achievable rate maximization based waveform design scheme, we propose a water-filling (WF) and sequential quadratic programming (SQP) based algorithm to derive the covariance matrix and the precoding matrix. Simulation results validate the proposed DFRC waveform designs and show that the achievable rate maximization based scheme has a better performance than the ISI minimization based scheme.
@@ I Introduction
The waveform design, which is one of the most challenging researches in the JRC system, is mainly classified into two categories. The first category is that the radar system and the communication system co-exist [10][11], where one system treats the other system as interference [12]. Therefore, multiplexing techniques including space division multiplexing (SDM) [13, 14, 15], time division multiplexing (TDM) [16, 17, 18], frequency division multiplexing (FDM) [19][20], and code division multiplexing (CDM) [21][22] have been applied to mitigate the interference in the co-existing system. However, in these multiplexing schemes, the co-existing waveform has a low utilization because the radar system and the communication system cannot employ the same resources, and two different kinds of waveforms are usually used respectively for radar and communication. The second category is that the radar system and the communication system are co-designed, which aims to develop a single waveform that can perform radar detection and communication functions simultaneously [23][24][25], and this waveform is called dual-functional radar-communication (DFRC) waveform. Compared with the co-existing waveform, the DFRC waveform has a higher degree of integration, thus it is more attractive to both academia and industry [26].
@@
All the above works adopted the single-carrier modulation, and the flat fading channel model was used. However, the frequency-selective multipath fading channel is more reasonable in the high-rate transmission. Orthogonal frequency division multiplexing (OFDM) [38][39] is one of the key technologies in the fifth generation (5G) mobile communication for its robustness to the multipath fading. Besides, it has been proven to be valid for the radar detection in [30]. Xu et al. in [40] presented a complete processing strategy for the measurement of range, velocity, and angle based on OFDM signals. Meanwhile, a trade-off scheme that employed dedicated subcarriers to balance the radar performance and the communication performance was also proposed. However, different subcarriers in [40] used the same precoding matrix, which led to a performance degradation in the DFRC system. The beam strength radiated towards the detection target is usually greater than that towards the communication user. Therefore, a subcarrier power allocation scheme was proposed in [41] to satisfy the different power requirements of the radar system and the communication system. However, [41] focuses on achieving the omnidirectional radiation in the radar detection phase, and it does not consider how to minimize CRB. Hu et al. in [42] extended the work of [23] and the PAPR of the designed OFDM waveform was optimized to achieve a better communication performance.
In this paper, we consider a single communication user MIMO system in the frequency-selective multipath fading environment. The DFRC base station (BS) utilizes a single waveform to realize the downlink communication and the target detection simultaneously. Because the MUI minimization scheme used in [23] cannot be directly used in the frequency-selective multipath fading environment, we propose an OFDM based waveform design scheme, which can effectively combat the multipath fading, to minimize the inter-stream interference (ISI) of the communication system and the CRB of the radar system. Furthermore, the achievable rate maximization based waveform design scheme and the joint water-filling and sequential quadratic programming (WF-SQP) algorithm are proposed to further improve the communication performance through the cooperation of different antennas.
@@
1. We focus on the performance trade-off between the communication system and the radar system, and propose two OFDM based DFRC waveform design schemes to combat the frequency-selective multipath fading. In the first proposed DFRC waveform design scheme, we minimize the ISI of the communication system and the CRB of the radar system. The algorithms for designing the transmit covariance matrix and the precoding matrix design algorithms are proposed.
2. Different from the first DFRC waveform design scheme, the achievable rate of the communication system is employed in the second DFRC waveform design scheme. The performance trade-off is also considered. To derive the optimal transmit covariance matrix, a joint water-filling and SQP algorithm is proposed. Furthermore, the precoding matrix is derived analytically.
3. Simulation results are provided and analyzed in detail. Simulation results show that the proposed DFRC waveform design schemes can achieve satisfied communication and radar performances. The second proposed DFRC waveform design scheme achieves better radar and communication performances than the first one.
The rest of this paper is organized as follows. Section II presents the MIMO-OFDM based system model. In Section III, we propose the first OFDM based DFRC waveform design scheme with the aim of minimizing the ISI of the communication and the CRB of the radar. In Section IV, we maximize the achievable rate of the communication system, and propose the second waveform design scheme. Simulation results are provided in Section V. Finally, the conclusion is drawn in Section VI.
@@ II System model
The transmit signal vector $\boldsymbol{x}(n)$ is expressed as
$\boldsymbol{x}(n)={[{x_{1}}(n),{x_{2}}(n),\cdots,{x_{{N_{T}}}}(n)]^{T}},n=1,2,\ldots,{N-1},$ (1)
where $N$ is the size of inverse discrete Fourier transform (IDFT), $x_{i}(n)$ is the time-domain OFDM signal transmitted through the $i$th antenna, and it is given by
@@
The frequency-domain signal vector $\boldsymbol{x}_{f}(k)$ and the time-domain signal vector $\boldsymbol{x}(n)$ are expressed as
$\boldsymbol{x}_{f}(k)={\left[{{X_{1}}(k),{X_{2}}(k),\cdots,{X_{{N_{T}}}}(k)}\right]^{T}},$ $k=0,1,\ldots,N-1,$ (3a)
$\boldsymbol{x}_{f}(k)=\boldsymbol{W}(k)$ $\boldsymbol{s}(k),$ (3b)
@@
$\boldsymbol{R}(i,s)$ $={1\over{{N^{2}}}}E\left\{{\sum\limits_{{k_{1}}=0}^{N-1}{\sum\limits_{{k_{2}}=0}^{N-1}{{X_{i}}({k_{1}})}}X_{s}^{*}({k_{2}}){e^{j{{2\pi}\over N}n({k_{1}}-{k_{2}})}}}\right\},$ (4b)
where $E\{\cdot\}$ denotes the expectation of a random variable, $(\cdot)^{H}$ is the conjugate transpose operator, $(\cdot)^{*}$ is the conjugate operator and $\boldsymbol{R}(i,s)$ is the $(i,s)$th entry of $\boldsymbol{R}$. It is supposed that each antenna employs all the subcarriers, and $E\{{{X}_{i}}({k_{1}}){X}_{s}^{*}({k_{2}})\}=0$ holds when $k_{1}\neq k_{2}$. Therefore, Eq. (4a) can be rewritten as
@@
$\boldsymbol{R}_{f}(k)=E\{\boldsymbol{x}_{f}(k){\boldsymbol{x}_{f}^{*}}(k)\}=\boldsymbol{W}(k){\boldsymbol{P}_{S}}(k){\boldsymbol{W}^{H}}(k).$ (6)
@@ II-A MIMO-OFDM Communication Model
${y_{s}}(n)=\sum\limits_{i=0}^{{N_{T}}-1}{{h_{s,i}}(n)\otimes{x_{i}}(n)}+{z_{s}}(n),s=1,2,\ldots,{N_{R}},$ (8)
where $\otimes$ denotes the convolution operation, and ${z_{s}}(n)$ is the noise component. Inter-subcarrier interference will increase the symbol error rate (SER) of the communication subsystem, and it will also cause inaccurate parameter estimation of the radar subsystem. It is assumed that ideal time-frequency synchronization is achieved at the receiver. In the frequency-domain, the received signal vector at the $k$th subcarrier can be expressed as
$\boldsymbol{Y}$ $(k)=\boldsymbol{H}(k)\boldsymbol{x}_{f}(k)+\boldsymbol{Z}(k)$ (9)
@@
where ${\boldsymbol{H}}(k)\in{\mathbb{C}^{{N_{R}}\times{N_{T}}}}$ is the frequency-domain channel matrix corresponding to the $k$th subcarrier, ${\boldsymbol{Z}}(k)\in{\mathbb{C}^{N_{R}}}$ is the frequency-domain Guassian noise vector with covariance matrix $\sigma_{\boldsymbol{Z}}^{2}{{\boldsymbol{I}}_{{N_{R}}}}$ , and ${\boldsymbol{I}}_{{N_{R}}}$ is an identity matrix with rank $N_{R}$. It is assumed that the base station utilizes the feedback information from communication users to obtain $\boldsymbol{H}(k)$.
@@
${C_{s}}=\gamma\sum\limits_{k=0}^{N-1}\sum\limits_{i=0}^{{N_{R}-1}}{{{\Delta f}\over B}}{{\log}_{2}}\left(\!{1\!+\!{{E\{{{{\left\|{{\boldsymbol{s}_{i}}(k)}\right\|}^{2}}}\}}\over{E\{{{{\left\|{{\boldsymbol{H}_{i}}(k)\boldsymbol{W}(k)\boldsymbol{s}(k)-{\boldsymbol{s}_{i}}(k)}\right\|}^{2}}}\}+\sigma_{\boldsymbol{Z}}^{2}}}}\!\right),$ (12)
where $\gamma={N\over{N+{N_{\rm{CP}}}}}$, $N_{\rm{CP}}$ denotes the length of OFDM cyclic prefix, $\Delta f$} is the subcarrier spacing, $B$ is the bandwidth of the OFDM system, ${\boldsymbol{s}_{i}}(k)$ is the $i$th entry of ${\boldsymbol{s}}(k)$, and ${\boldsymbol{H}_{i}}(k)$ is the $i$th row of ${\boldsymbol{H}}(k)$. Note that the effect of the cyclic prefix to the data rate is ignored.
@@
${C_{t}}=\gamma\cdot{\rm{max}}\bigg{\{}\sum\limits_{k=0}^{N-1}{{{\Delta f}\over B}}{{\log}_{2}}\left({\det\left({{\boldsymbol{I}_{{N_{R}}}}+{{\boldsymbol{H}(k){\boldsymbol{R}_{f}}(k){\boldsymbol{H}^{H}}(k)}\over{\sigma_{\boldsymbol{Z}}^{2}}}}\right)}\right)\bigg{\}},$ (13)
@@
Besides, the proposed scheme can be applied in the multiple communication user scenario, and time division multiple access (TDMA) or orthogonal frequency division multiple access (OFDMA) techniques can be used.
@@ II-B MIMO Radar Model
where $\zeta$ is the reflection coefficient and contains the information of the radar cross section (RCS) of the target, ${\boldsymbol{z}_{r}}(n)$ is the Gaussian noise vector term which includes multipath interference from the same target, $\boldsymbol{A}(\theta)=\boldsymbol{a}(\theta)\boldsymbol{a}(\theta)^{H}$, $\theta$ is the angle of arrival of the target, and the steering vector $\boldsymbol{a}(\theta)$ is given by [35]
@@
where the center of the ULA antennas is chosen as the reference point, and $\Delta$ is the normalized antenna spacing. It should be noted that the normalized steering vector for different subcarriers may have slight differences, which are neglected in this paper because the frequency difference between subcarriers is much less than the carrier frequency.
CRB represents the lower bound of the variance of an unbiased estimator, and is an important indicator of the radar performance. In this paper, a smaller CRB represents a stronger stability of the radar parameter estimation. The CRB of $\theta$ can be derived as [44]
@@
where $\boldsymbol{\dot{A}}(\theta)={{\partial\boldsymbol{A}(\theta)}\over{\partial\theta}}$, and $\rm{tr}(\cdot)$ is the operator of matrix trace. Substituding $\boldsymbol{A}(\theta)=\boldsymbol{a}(\theta)\boldsymbol{a}(\theta)^{H}$ into Eq. (16), the CRB of $\theta$ can be rewritten as [44]
@@ III Joint waveform design based on CRB and ISI minimization
In this section, we first design the time-domain covariance matrix to minimize the CRB of $\theta$. After that, an optimization problem of minimizing the interference among different information streams for a given time-domain covariance matrix is formulated, and the frequency-domain covariance matrix and the corresponding precoding matrix are designed. Besides, we propose the joint waveform design scheme based on the CRB and ISI minimization to achieve the performance trade-off.
@@ III-A Radar-only problem formulation
$\rm{OP1:}\quad\mathop{\arg\min}\limits_{\boldsymbol{R}}$ ${1\over{2\rm{SNR}{{\left\|{\boldsymbol{\dot{a}}(\theta)}\right\|}^{2}}\boldsymbol{a}{{(\theta)}^{H}}\boldsymbol{R}\boldsymbol{a}(\theta)}}$ (18a)
@@
where $P_{T}$} is the total transmit power. The optimization problem OP1 is convex [45] and has a closed-form solution. The optimal time-domain covariance matrix is given by
$\boldsymbol{R}_{d}={{{P_{T}}}\over{{N_{T}}}}{\boldsymbol{a}^{*}}(\theta){\boldsymbol{a}^{T}}(\theta).$ (19)
Although we focus on the single detection target scenario in this paper, the proposed scheme can also be extended to the multi-target scenario.